1:Stock R has a beta of 2.5, Stock S has a beta of 0.95, the required return on an average stock is 11%, and the risk-free rate of return is 3%. By how much does the required return on the riskier stock exceed the required return on the less risky stock? Round your answer to two decimal places. %

2:Beale Manufacturing Company has a beta of 2, and Foley Industries has a beta of 0.35. The required return on an index fund that holds the entire stock market is 12%. The risk-free rate of interest is 3.25%. By how much does Beale's required return exceed Foley's required return? Round your answer to two decimal places. %

3: You have been managing a $5 million portfolio that has a beta of 0.70 and a required rate of return of 12%. The current risk-free rate is 6.25%. Assume that you receive another $500,000. If you invest the money in a stock with a beta of 1.45, what will be the required return on your $5.5 million portfolio? Do not round intermediate calculations. Round your answer to two decimal places.

%
Answer #1

1.Stock R is more risky than stock S because of its beta which is higher.
The required return on the riskier stock exceed the required return on the less risky stock=(Beta Stock R-Beta Stock S)*(Required Rate of Market-Risk free rate) =(2.5-0.95)*(11%-3%) =12.4%

2. Beale's required return exceed Foley's required return =(Beta Beagle-Beta Foley )*(Required Rate of Market-Risk free rate)
=(2-0.35)*(12%-3.25%) =14.44%